The electric potential at x=3.00 m is 120 V, and the electric potential at x=5.00 m is 190 V. What is the x-component of the electric in this region, assuming the field is uniform?
1
Expert's answer
2020-11-10T09:56:56-0500

The x component of electric field is

"E_x=-\frac{dV}{dx}=--(\frac{190-120}{5.00-3.00}) = - 35 V"